OpenStudy (firejay5):

I don't know how to do it, so you need to help me out here. Solve each system of equations, and show your work/steps: 12. 2x - y = 2 3z = 21 4x + z = 19 14. 5x + 2y = 4 3x + 4y + 2z = 6 7x + 3y + 4z = 29

OpenStudy (anonymous):

12. Start off by solving for z 3z=21 z=7

OpenStudy (anonymous):

You can use this z value to plug into the equation 4x+z=19 in order to solve for x 4x+7=19 4x=12 x=3

OpenStudy (anonymous):

Finally, plug in your x value into the equation 2x-y=2 in order to solve for y 2(3)-y=2 6-y=2 y=4

OpenStudy (anonymous):

This method is known as the substitution method and is one way of solving systems of equations

OpenStudy (firejay5):

did you subtract 2 or 6

OpenStudy (anonymous):

I subtracted 6 from 2 -y=-4 negatives cancel out y=4 But subtracting 2 from 6 would work as well 4-y=0 y=4

OpenStudy (firejay5):

wouldn't you get y = -4, if I subtracted 2

OpenStudy (anonymous):

You would subtract 2 from 6, and be left with 4-y=0 Add the y to the other side → y=+4

OpenStudy (firejay5):

@Copythat What would I do here

OpenStudy (anonymous):

The easiest method for solving 14. is elimination

OpenStudy (anonymous):

14. 5x + 2y = 4 3x + 4y + 2z = 6 7x + 3y + 4z = 29 [5x+2y=4]*3 → 15x+6y=12 [3x+4y+2z=6]*4 → 15x+16y+8z=24 15x+6y=12 -15x+16y+8z=24 ____________________ -10y+8z=-12

OpenStudy (anonymous):

I took the first two equations and multiplied the first by the x value in the second and the second by the x value in the first. Your goal here is to eliminate the x variable.

OpenStudy (firejay5):

(3x + 4y + 2z = 6)*-5

OpenStudy (anonymous):

[3x+4y+2z=6]*5 → 15x+20y+10z+30 Once you have both of your new equations, you subtract in order to eliminate your x variable

OpenStudy (firejay5):

@Copythat understand

OpenStudy (anonymous):

Its not a mistake. Subtracting 15 from 15 gives you 0.

OpenStudy (anonymous):

If you multiply by -5, then you must use addition to eliminate x

OpenStudy (firejay5):

no you don't

OpenStudy (anonymous):

Why? You now have a new equation which you can use to plug into the third equation and solve for another variable

OpenStudy (firejay5):

(5x + 2y = 4)*3 ----> 15x + 6y = 12 (3x + 4y + 2z)*-5 =6 --> -15x - 20y - 10z = -30 15x + 6y = 12 -15x - 20y -10z = -30 0x - 14y - 10z = -18

OpenStudy (firejay5):

@ajprincess What do I do now???

OpenStudy (ajprincess):

Nw take the two eqns 5x + 2y = 4 and 7x + 3y + 4z = 29. Eliminate x terms from both eqns. Can u do that @Firejay5?

OpenStudy (firejay5):

Yea I'll give it a shot and come back with an answer

OpenStudy (ajprincess):

ha k. 5n:)

OpenStudy (firejay5):

35x + 14y = 28 -35x -15y - 20z = -130 -y - 20z = - 102

OpenStudy (firejay5):

is that correct @ajprincess

OpenStudy (ajprincess):

29*-5=-145

OpenStudy (ajprincess):

35x + 14y = 28 -35x -15y - 20z = -145 -y - 20z = - 117

OpenStudy (ajprincess):

- 14y - 10z = -18 -y - 20z = - 117 Nw try eliminating y from both the eqns. Can u do it @FireJay5?

OpenStudy (firejay5):

yes

OpenStudy (ajprincess):

after doing it solve for z. k?

OpenStudy (firejay5):

you should've gotten -130, not -145

OpenStudy (ajprincess):

 What is 29*-5=?